OpenStudy (anonymous):

What is the correct reason for statement 5? Prove: 5y + 4 + 6(1 + y) = 11y + 10 Statements Reasons 1. 5y + 4 + 6(1 + y) 1. Given 2. 5y + 4 + (6+ 6y) 2. Distributive Property 3. 5y + (4 + 6) + 6y 3. Associative Property 4. 5y + 10 + 6y 4. Addition 5. 5y + 6y + 10 5. 6. 11y + 10 6. Addition A. Multiplication B. Associative Property C. Commutative Property D. Addition @IMStuck

OpenStudy (anonymous):

i think it is c?

OpenStudy (anonymous):

what did you do going from 4 to 5?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

nope

OpenStudy (texaschic101):

you rearranged your numbers

OpenStudy (anonymous):

re arranged so it's addition?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sorry i wasnt paying close attention @pgpilot326 can you help me with one more?

OpenStudy (texaschic101):

look at this.. http://www.mathsisfun.com/associative-commutative-distributive.html

OpenStudy (anonymous):

you moved them just like a person moves from their to work (each day they work). that's called a commute. so what property do you think you use to move numbers around like that?

OpenStudy (anonymous):

home to work

OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 5x + 9 • x • 3 = 32x 1. 5x + 9 • x • 3 = 5x + 9 • 3 • x 1. ? 2. 5x + 9 • x • 3 = 5x + 27x 2. Multiplication 3. 5x + 9 • x • 3 = 32x 3. ? A. Associative Property of Addition, Addition B. Associative Property of Multiplication, Addition C. Commutative Property of Multiplication, Addition D. Commutative Property of Multiplication, Multiplication I think it is b?

OpenStudy (noahbred):

wait yo dude is this k12